# Towards Cross-Lingual LLM Evaluation for European Languages

### v1

@@ 4.2 Impact of available training corpora
To better analyze the performance of LLMs across different languages, we categorized our target languages into high-resource (HRL) and medium-resource (MRL) groups, based on their proportion in the CommonCrawl dataset, which serves as the primary pre-training source for many LLMs. High-resource languages (HRL) had a CommonCrawl representation of 1% or more, whereas medium-resource languages (MRL) had between 0.01% and 1%. The language statistics accompanied with evaluation results of the LLama-3.1-70B-Instruct Dubey et al. (2024) model is depicted in Table 1.
As can be seen there is a downwards trend with outliers in Swedish and Danish, both Germanic languages that might profit from English or other Germanic training corpora (cf. Table 1).
@@ 4.3 Influence of Language Family
Similarly, the Meta-Llama-3.1-70B-Instruct Dubey et al. (2024) model achieves average accuracies of 0.745 (Germanic), 0.726 (Romance), and 0.680 (Slavic), reinforcing the trend of lower performance on Slavic languages. In the following we compare the model performance on similar sized (in CC %) languages such as Romanian (Romance, 0.52%), Swedish (Germanic, 0.63%) and Polish (Slavic, 1.09%). Table 1 indicates that even though, Swedish is less represented than Polish, it achieves higher performance across all tasks. Also Romanian as the smallest represented language achieves higher performance across all five tasks, indicating that the language family is of influence in the performance analysis.
5 Correlation with Human Preference Scores
We examine the correlation between our introduced EU21 benchmarks and the Elo scores from the LMSYS Chatbot Arena [^16] https://chat.lmsys.org/?leaderboard to better understand how human preferences are reflected.
The Elo scores we use are derived from a dataset [^17] https://storage.googleapis.com/arena_external_data/public/clean_battle_20240814_public.json of about 1.8 million human quality judgements. In the LMSYS Chatbot Arena Chiang et al. (2024) users can vote on a pair of LLM responses in anonymized and randomized "battles".
The battles are divided into several categories, some of which correspond to the response languages. The Elo rankings are computed independently for each category.
In order to quantify how well the accuracies of a given model on our translated datasets correlate with the Elo scores, we compute the Pearson and Spearman Correlation Coefficients between the accuracy and the Elo scores of the $n=17$ different models. We compute these correlation separately for the Czech, French, German, Italian, Polish, Portuegese, Spanish and for English as a reference point. The remaining mostly lower ressource languages have a high uncertainty on their Elo scores as fewer battles (cf. Table 20) are available.
The LLMs for the analysis were mostly chosen based on the Elo rankings available, according to the following criteria:
@@ 5 Correlation with Human Preference Scores
  • LMSYS Elo Ratings available for English, French, German and Spanish
@@ 6 Influence of the translation service
While OKAPI Lai et al. (2023) and our evaluation datasets share a similar structure, they differ in key aspects such as translation methodology and sample alignment. The most notable difference between the two datasets is the translation approach: The OGX dataset was translated using DeepL, whereas OKAPI employed ChatGPT for this task. The translations for the OKAPI tasks cover 31 languages, including both EU and non-EU languages. For the OpenGPT-X evaluation datasets, we focused on translations into 20 European languages. The common and new set of EU languages between the OKAPI and OpenGPT-X are shown in Table 3.
@@
Table 3: Comparison of languages between OKAPI and OpenGPT-X Datasets
In this section, we present a correlation analysis with a focus on the ARC, HellaSwag, and MMLU tasks, alongside a COMET score Rei et al. (2023) evaluation to assess the translation quality across both datasets.
The idea behind the correlation analysis is to record [^18] The evaluation was performed using the LM-Evaluation-Harness framework, with sample-level logging enabled.whether the model responded correctly or incorrectly for each OKAPI and OGX sample and to examine if the predictions are correlated, allowing direct performance comparison on sample-level on the same data. For a sample-level comparison, we ensured that OGX and OKAPI samples are matched and can be linked to their original translations. While ARC, HellaSwag, and MMLU samples have IDs matching the original English subsets, enabling mapping, this does not apply to TruthfulQA.
For the correlation analysis, we calculated the correlation matrices for different models, tasks, and languages, listed in Appendix A.
@@
As seen in Table 19, ARC has the highest correlation at 0.61, followed by Hellaswag and MMLU (both 0.54). ARC’s structure and reliance on factual knowledge likely make model performance more stable and predictable.
Regarding the task performance correlation, we can see from Table 18 that correlations range from 0.44 for Hungarian to 0.59 for Spanish. Hungarian’s lower score may reflect fewer training resources, while Spanish likely benefits from better data availability, leading to more consistent model performance.
Translation quality likely plays a role in these results, as ambiguous or incorrect translations, poorly adapted prompts, or task mismatches can negatively affect model performance. Thus, the observed correlations may also reflect the degree to which translation quality impacts each language’s evaluation results. Table 5, Table 6, Fig. 4 and Table 7 The examples in Appendix A Table 5, Table 6, Fig. 4 and Table 7 underscore the importance of context-aware translation in multilingual tasks, as naive translations can reduce clarity and coherence, ultimately affecting model performance.
As part of the further analysis, we determined the contingency table (4) to compare the answers of the models on our evaluation datasets and the OKAPI datasets. Both evaluations show similar alignment, with agreements on correct answers (38.1%) and incorrect answers (39.3%) being fairly balanced. However, OKAPI seems to perform slightly better in cases of disagreement, as OKAPI is correct and our model is incorrect in 11.8% of cases, compared to 10.8% where our model is correct and OKAPI is incorrect. This suggests that, while the performance of both models is comparable, OKAPI might have a slight edge in terms of accuracy in cases where they disagree.
@@
Further we used the COMET score Rei et al. (2023) to analyze the translation quality of both datasets. Since there is no target translation for the different benchmarks, we evaluated the translation quality of the OGX and OKAPI datasets using the COMET-KIWI metric Rei et al. (2023). This metric uses a pre-trained language model to predict translation quality. A score of 1 indicates a high-quality translation and a score close to 0 indicates a bad translation. Both the OGX and OKAPI datasets have a relatively high average score of 0.8208 and 0.8190, respectively, across all tasks and translations. The exact scores can be found in the Appendix in Table Table 29. It is important to highlight that because there is no target translation, no statistical significance can be calculated between the scores of OKAPI and OGX. Therefore, we cannot conclude that the OGX translations are automatically higher quality based on the COMET KIWI score. Nevertheless, it can be shown that the OGX datasets have a consistently high translation quality across all languages (See Appendix in Table Table 29).
We additionally compared all our models on the ARC benchmark Fig. 10, Hellaswag Fig. 11 and MMLU Fig. 12. Overall we can see that all evaluated models achieve higher accuracies on our tasks compared to OKAPI. This analysis highlights the extent to which our and OKAPI results align across models, tasks, and languages, providing insights into the consistency and reliability of model evaluations across different datasets.
@@ 8 Limitations
A hindrance in achieving benchmarks comparable across languages is localization of cultural aspects. Toxicity benchmarks in particular can suffer from this, as stereotypes and terms related to them likely do not translate directly been cultures and hence languages. To a lesser extent this applies to all benchmarks requiring some degree of common sense and world knowledge. A team of reseachers has launched an initiative [^19] https://huggingface.co/spaces/CohereForAI/MMLU-evaluationto correct some of these cultural biases in multilingual datasets (predominantly MMLU), through crowd-sourced manual work. While we are aware of this initiative, we opted not to conduct a similar enterprise in order to avoid doubled efforts. Aligning the results from the initative with our own translations will be an avenue for further work.
@@
Furthermore the degree to which a translation for an given task is not always feasible. The pronoun-resolution benchmark WinoGrande Sakaguchi et al. (2019) for instance is posed as a multiple choice task, where a on object or subject in a sentence is masked and each choice constitutes a replacement of the mask. The model is then tasked with selecting the sentence that is semantically correct. Tables LABEL:fig:winogrande_base and LABEL:fig:winogrande_spa demonstrate how in languages with gendered nouns or gendered adjectives in predicative use, the benchmark becomes easier due to strong grammatical signals indicating the correct choice of subject or object. It is for that reason we decided against translation of WinoGrande.
In order to compare the translations of OGX with the original benchmark or other translations such as OKAPI, the samples must be mapped. This requires a one-to-one mapping. However, this mapping was impossible for all benchmarks; therefore, we could only compare three benchmarks with the OKAPI dataset: Hellaswag, MMLU, and ARC. For Hellaswag, we used the ind and the source_id; for ARC, we used the id; and for MMLU, the subtask_name and the natural sorting of the benchmarks, which turned out to be a one-to-one mapping.
We analyzed the translated sequences with unexpectedly low scores (<0.5) during the quality analysis of the translations using the COMET KIWI score. Depending on the benchmark, some sequences were not translated. As a result, some answer options are either in the original language or do not appear at all. The maximum error rate is 2.3 % for Hellaswag, 1.8 % for ARC, and 2.1 % for MMLU across all languages. Although the errors show no significant difference in direct comparison with OKAPI (see Section 6) nor with the original benchmarks, the analysis using the COMET score in the translation pipeline should be considered in future work.
No measures to indicate overfitting on English benchmarks, such as math-based benchmarks as in https://arxiv.org/pdf/2405.00332
@@ 1. Kampfroboter des Bösen und endet mit einem "to be continued"
Appendix B European LLM Evaluation Results

### v2

@@ 4.2 Impact of available training corpora
To better analyze the performance of LLMs across different languages, we categorized our target languages into high-resource (HRL) and medium-resource (MRL) groups, based on their proportion in the CommonCrawl dataset, which serves as the primary pre-training source for many LLMs. High-resource languages (HRL) have a CommonCrawl representation of 1% or more, whereas medium-resource languages (MRL) have between 0.01% and 1%. The language statistics accompanied with evaluation results of the LLama-3.1-70B-Instruct Dubey et al. (2024) model is depicted in Table 1.
There is a downwards trend with outliers in Swedish and Danish, both Germanic languages that might profit from English or other Germanic training corpora (cf. Table 1). The performance of EU21-TQA, EU21-MMLU, and EU21-GSM8K shows a slight decline in correlation with the number of speakers in millions and the percentage share of CommonCrawl data across different languages, while EU21-ARC and EU21-Hellaswag exhibit a more pronounced performance drop.
Both EU21-ARC and EU21-Hellaswag are benchmarks that focus on reasoning, commonsense knowledge, and context understanding. These tasks might require more language specific knowledge, which may not be as well represented in languages with fewer speakers or less CommonCrawl data. In contrast, tasks like MMLU and GSM8K, while challenging, may rely more on factual recall, basic math, or language understanding. We hypothesize these are capabilities a model can easier transfer across languages.
EU21- EU21- EU21- EU21- EU21-
@@ 4.3 Influence of Language Family
Similarly, the Meta-Llama-3.1-70B-Instruct Dubey et al. (2024) model achieves average accuracies of 0.745 (Germanic), 0.726 (Romance), and 0.680 (Slavic), reinforcing the trend of lower performance on Slavic languages.
In the following we compare the model performance on similar sized (in CC %) languages such as Romanian (Romance, 0.52%), Swedish (Germanic, 0.63%) and Polish (Slavic, 1.09%). Table 1 indicates that even though, Swedish is less represented than Polish, it achieves higher performance across all tasks. Also Romanian as the smallest represented language achieves higher performance across all five tasks, indicating that the language family is of influence in the performance analysis.
5 Correlation with Human Preferences
We examine the correlation between the introduced EU21 benchmarks and the Elo scores from the LMSYS Chatbot Arena [^16] https://chat.lmsys.org/?leaderboard to better understand how human preferences are reflected.
In the LMSYS Chatbot Arena Chiang et al. (2024), users can vote on a pair of LLM responses in anonymized and randomized "battles". The resulting preference data are used to compute Elo rankings published on the LMSYS chatbot arena leaderboard. There are multiple independently computed rankings in different categories, some of which correspond to the languages of the conversations. We utilize an input dataset with timestamp 20240814 containing about 1.8 million human quality judgements provided for reproducing the rankings (cf. Section C.1). With these data, we compute the Elo scores using FastChat171717commit e5dc446f.
The LLMs for the subsequent correlation analysis were mostly chosen using the following criteria:
@@ 5 Correlation with Human Preferences
To quantify how well the accuracies of a given model on our translated datasets correlate with the Elo scores, Then we compute the Pearson and Spearman Correlation Coefficients between the accuracy and the Elo scores of $n=17$ models using scipy.stats with default settings and confidence level 95%. We compute these correlations for Czech, French, German, Italian, Polish, Portuguese, Spanish, as for these languages, at least 50% of the Elo ranks of the models we consider can be distinguished with about 95% confidence, and for English as a reference point. The remaining mostly smaller medium-resource languages have a high uncertainty in their Elo scores as fewer battles (cf. Table 21) are available.
@@ 6 Influence of the translation service
While OKAPI Lai et al. (2023) and our evaluation datasets share a similar structure, they differ in key aspects such as translation methodology and sample alignment. The most notable difference between the two datasets is the translation approach: Our dataset was translated using DeepL, whereas OKAPI employed ChatGPT for this task. The translations for the OKAPI tasks cover 31 languages, including both EU and non-EU languages. For our evaluation datasets, we focused on translations into 20 European languages. The common and new set of EU languages between the OKAPI and our benchmarks are shown in Table 3.
@@
Table 3: Comparison of languages between OKAPI and our Datasets
In this section, we present a correlation analysis with a focus on the EU20-ARC, EU20-HellaSwag, and EU20-MMLU tasks, alongside a COMET score Rei et al. (2023) evaluation to assess the translation quality across both datasets.
The idea behind the correlation analysis is to record [^18] The evaluation was performed using the LM-Evaluation-Harness framework, with sample-level logging enabled.whether the model responded correctly or incorrectly for each OKAPI sample and sample from our dataset and to examine if the predictions are correlated, allowing direct performance comparison on sample-level on the same data. For a sample-level comparison, we ensured that our examples and OKAPI samples are matched and can be linked to their original translations. While EU20-ARC, EU20-HellaSwag, and EU20-MMLU samples have IDs matching the original English subsets, enabling mapping, this does not apply to EU20-TruthfulQA.
For the correlation analysis, we calculated the correlation matrices for different models, tasks, and languages, listed in Section D.3.
@@
As seen in Table 32, RC has the highest correlation at 0.61, followed by HellaSwag and MMLU (both 0.54). ARC’s structure and reliance on factual knowledge likely make model performance more stable and predictable.
Table 31 shows that correlations range from 0.44 for Hungarian to 0.59 for Spanish. Hungarian’s lower score may reflect fewer training resources, while Spanish likely benefits from better data availability, leading to more consistent performance.
Translation quality likely plays a role in these results, as ambiguous or incorrect translations, poorly adapted prompts, or task mismatches can negatively affect model performance. Thus, the observed correlations may also reflect the degree to which translation quality impacts each language’s evaluation results. Examples illustrating translation quality issues in the ten common languages can be found in LABEL:tab:EU21_okapi_hellaswag for the HellaSwag task.
As part of the further analysis, we determined the contingency table (Table 4) to compare the answers of the models on our evaluation datasets and the OKAPI datasets. Both evaluations show similar alignment, with agreements on correct answers (38.1%) and incorrect answers (39.3%) being fairly balanced. However, OKAPI seems to perform slightly better in cases of disagreement, as OKAPI is correct and our model is incorrect in 11.8% of cases, compared to 10.8% where our model is correct and OKAPI is incorrect. This suggests that, while the performance of both models is comparable, OKAPI might have a slight edge in terms of accuracy in cases where they disagree.
@@
Further we used the COMET score Rei et al. (2023) to analyze the translation quality of both datasets. Since there is no target translation for the different benchmarks, we evaluated the translation quality of the our and OKAPI datasets using the COMET-KIWI metric Rei et al. (2023). This metric uses a pre-trained language model to predict translation quality. A score of 1 indicates a high-quality translation and a score close to 0 indicates a bad translation. Both our and OKAPI datasets have a relatively high average score of 0.8208 and 0.8190, respectively, across all tasks and translations. The exact scores can be found in the Appendix in Table Table 33. It is important to highlight that because there is no target translation, no statistical significance can be calculated between the scores of OKAPI and ours. Therefore, we cannot conclude that our translations are automatically higher quality based on the COMET KIWI score. Nevertheless, it can be shown that our datasets have a consistently high translation quality across all languages (cf. Table Table 33).
We additionally compared all our LLMs on the ARC, (Figure 10), HellaSwag (Figure 11) and MMLU benchmarks (Figure 12).
Overall we can see that all evaluated LLMs achieve higher accuracies on our tasks compared to OKAPI. This analysis highlights the extent to which our and OKAPI results align across models, tasks, and languages, providing insights into the consistency and reliability of model evaluations across datasets.
@@ 8 Limitations
It depends much on the task whether translations are a good way to create corpora in parallel languages. The ambiguity resolution benchmark WinoGrande Sakaguchi et al. (2019) for instance is posed as a multiple choice task, where a object or subject in a sentence is masked and each choice constitutes a replacement of the mask. The model is then tasked with selecting the sentence that is semantically correct. Table 5 and 6 in the Appendix demonstrate how in languages with gendered nouns or gendered adjectives in predicative use, the benchmark becomes easier due to strong grammatical signals indicating the correct choice of subject or object. Depending on the benchmark, some sequences were sporadically not translated, so some answer options are either in the original language (English) or are left blank. We detected the issue by analyzing the translated sequences with unexpectedly low COMET KIWI scores (<0.5) during the quality analysis of the translations. The maximum error rate for any language (assuming no errors for scores $\geq 0.5$) is 2.3% for HellaSwag (IT), 1.8% for ARC (ET), and 2.1% for MMLU (PT) across all languages.
Variations in accuracy were noted when testing language models with different versions of the LM Eval Harness framework (v0.4.1, v0.4.3, and 42dc24/main). The differences (less than 1%), affected ARC and MMLU, were attributed mainly to changes in batch size, and task formatting. These variations emphasize the importance of maintaining consistent evaluation practices to ensure accurate model comparisons across different versions. Details can be found in section B.2.
9 Ethical & Broader Impact
The ability to evaluate large language models (LLMs) across a wide range of European languages, particularly underrepresented ones, is a critical step toward enhancing inclusivity and accessibility in natural language processing (NLP). By ensuring that LLMs can perform well in languages beyond English or other high-resource languages, we contribute to a more equitable digital landscape where speakers of less widely spoken languages have equal access to advanced language technologies. However, this inclusivity brings unique challenges, particularly in achieving benchmarks that are comparable across diverse linguistic and cultural contexts.
Acknowledgments
This work was funded by the Federal Ministry of Education and Research of Germany and the state of North-Rhine Westphalia as part of the Lamarr-Institute for Machine Learning and Artificial Intelligence, LAMARR22B as well as by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) through the project OpenGPT-X (project no. 68GX21007D) and by the European Union’s Horizon 2020 research and innovation program under grant agreement No 101135671 (TrustLLM). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC) as well as the Center for Information Services and High Performance Computing [Zentrum für Informationsdienste und Hochleistungsrechnen (ZIH)] at TU Dresden for providing its facilities for automatic evaluation computations.
@@ Appendix A Experimental Setup
A.1 Problematic Examples
Tables 5 and 6 contain a sample from the development split of WinoGrande-XL and its translation, illustrating challenges in 1:1 translation of evaluation datasets.
Sentence: The juice from the mango fruit could not fill up the cup because the _ is small.
Choices:
mango
cup
Table 5: Development split instance of the WinoGrande-XL dataset.
Choices:
El zumo de la fruta de mango no podía llenar la taza porque el mango es pequeño.
El zumo de la fruta de mango no podía llenar la taza porque la taza es pequeña.
Table 6: Potential Spanish translation of the instance in Table 5. Highlighted in red are differences between the two choices beyond replacement of the subject.
@@ 1. Kampfroboter des Bösen und endet mit einem "to be continued"
Appendix B Evaluation Results
B.1 Model overview
Table 10: Model Overview; Legend: D: Used in downstream evaluation (cf. Section 4); T: Used in OKAPI translation comparison (cf. Section 6); C: Used in LMSYS correlation analysis (cf. Section 5)
Model D T C Hugging Face Link (https://hf.co/...)
Aya-23-8B CohereForAI/aya-23-8B
Bloom-7b1 bigscience/bloom-7b1
Bloomz-7b1 bigscience/bloomz-7b1
c4ai-command-r-35B-v01 CohereForAI/c4ai-command-r-v01
EuroLLM-1.7B utter-project/EuroLLM-1.7B
EuroLLM-1.7B-Instruct utter-project/EuroLLM-1.7B-Instruct
Gemma-7b google/gemma-7b-it
Gemma-1.1-7b-Instruct google/gemma-1.1-7b-it
Gemma-2-9b-Instruct google/gemma-2-9b-it
Gemma-2-27b-Instruct google/gemma-2-27b-it
Meta-Llama-2-7B meta-llama/Llama-2-7b-hf
Meta-Llama-2-7B-Chat meta-llama/Llama-2-7b-chat-hf
Meta-Llama-2-13B-Chat meta-llama/Llama-2-13b-chat-hf
Meta-Llama-3-8B meta-llama/Meta-Llama-3-8B
Meta-Llama-3-8B-Instruct meta-llama/Meta-Llama-3-8B-Instruct
Meta-Llama-3.1-8B meta-llama/Llama-3.1-8B
Meta-Llama-3.1-8B-Instruct meta-llama/Llama-3.1-8B-Instruct
Meta-Llama-3.1-70B-Instruct meta-llama/Llama-3.1-70B-Instruct
Mistral-7B-v0.1 mistralai/Mistral-7B-v0.3
Mistral-7B-v0.3 mistralai/Mistral-7B-v0.3
Mistral-7B-Instruct-v0.1 mistralai/Mistral-7B-Instruct-v0.1
Mistral-7B-Instruct-v0.2 mistralai/Mistral-7B-Instruct-v0.2
Mistral-7B-Instruct-v0.3 mistralai/Mistral-7B-Instruct-v0.3
Mistral-Nemo-Base-12.2B_2407 mistralai/Mistral-Nemo-Instruct-2407
Mistral-Nemo-Instruct-12.2B_2407 mistralai/Mistral-Nemo-Instruct-2407
Mistral-NeMo-Minitron-8B-Base nvidia/Mistral-NeMo-Minitron-8B-Base
Mixtral-8x7B-v0.1 mistralai/Mixtral-8x7B-v0.1
Mixtral-8x7B-Instruct-v0.1 mistralai/Mixtral-8x7B-Instruct-v0.1
Occiglot-7b-eu5 occiglot/occiglot-7b-eu5
Occiglot-7b-eu5-Instruct occiglot/occiglot-7b-eu5-instruct
Pharia-1-LLM-7B-control Aleph-Alpha/Pharia-1-LLM-7B-control-hf
Pharia-1-LLM-7B-control-aligned Aleph-Alpha/Pharia-1-LLM-7B-control-aligned-hf
Phi-3-medium-14B-4k-Instruct microsoft/Phi-3-medium-4k-instruct
Phi-3-medium-14B-128k-Instruct microsoft/Phi-3-mini-128k-instruct
Phi-3-mini-4k-Instruct microsoft/Phi-3-mini-4k-instruct
Phi-3-mini-3.8B-128k-Instruct microsoft/Phi-3-mini-128k-instruct
Qwen2-7B Qwen/Qwen2-7B
Qwen2-7B-Instruct Qwen/Qwen2-7B-Instruct
Vicuna-13b-v1.5 lmsys/vicuna-13b-v1.5
Vicuna-33b-v1.3 lmsys/vicuna-33b-v1.3
Table LABEL:tab:eval_model_overview contains an overview of the models used in the different evaluations performed in this paper. The columns D, T and C indicate whether the respective model was used in section 4, 6 and 5, respectively. The URLs for to the Hugging Face repositories of the models had to be shortened due to page width constraints and have to be prefixed with https://hf.co.
B.2 Accuracy Variations
Notable accuracy variations were observed when evaluating the performance of language models (LMs) using different versions of the LM Eval Harness (v0.4.1, v0.4.3, and 42dc24/main). The evaluations described in the following focused on the French version of the ARC-challenge dataset, and the findings highlight key differences and potential causes of these discrepancies. The results were conducted using Mistral-7B-v0.1, and accuracy was assessed across different batch sizes and floating-point formats. The model evaluations used batch sizes of 30, 32, and 60. These are the key observations and the numbers that underline the findings:
Accuracy Variations Across Versions:
Small but consistent discrepancies in accuracies were noted across the different versions. For instance, the accuracies in v0.4.1 and v0.4.3 for a batch size of 30 were both 40.27% for the ARC challenge, while the normalized accuracy varied slightly, i.e. v0.4.1 46.59% and v0.4.3 46.41%. When using a larger batch size of 60, both versions again reported an accuracy of 40.27%, with a normalized accuracy of v0.4.1 46.50% and v0.4.3 46.59%
Base Version BS Acc_Norm Acc
v0.4.1 30 0.46587 0.40273
v0.4.1 32 0.46587 0.40188
v0.4.1 60 0.46502 0.40273
v0.4.3 30 0.46502 0.40273
v0.4.3 32 0.46416 0.40188
v0.4.3 60 0.46587 0.40273
42dc24 30 0.46502 0.40273
Table 11: Evaluation results with different LM Eval Harness base versions, batch sizes, and metrics.
Potential Causes for Discrepancies:
These differences suggest that changes in batch size and GPU counts influenced how padding samples were handled, causing minor variations in accuracy. This observation is consistent with the findings of other research groups using the LM Eval Harness for benchmarking  [^19] https://huggingface.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard. Differences in the usage of random number generators in few-shot settings were noted, particularly between v0.4.1 and v0.4.3, which could affect model accuracy. In zero-shot settings, the accuracy difference was minimal, and evaluating in a few-shot setting did not significantly impact the outcome. Based on a manual code inspection, we suspect that few-shot results may vary depending on what other tasks are evaluated in the same run. Furthermore, sample logs revealed that in data-parallel mode, the same sequence of few-shot contexts is used in each process, such that only the actual test sample differed between the $i$-th samples of each process. This might possibly increase the standard error depending on how many processes (GPUs) are used as statistical noise caused by the specific selection of few-shot contexts might be amplified. We also compared float16 and float32 runs, and results showed no significant differences in accuracy between these formats, e.g. float16 accuracy 40.53% (in both versions) and float32 accuracy 40.44% (in both versions).
Key Differences Identified:
A primary source of accuracy discrepancies was found to be the changed reordering and collation of batches in v0.4.3. The accuracy results for reverse-patched reordering matched the v0.4.1 results, confirming that the reordering was the cause in this setting. Specifically, the v0.4.3 reverse-patched reordering exhibited 40.27% Accuracy and a Normalized Accuracy of 46.59% (exactly matching v0.4.1) Task specification fields previously treated as plain texts were re-interpreted as Jinja2 templates in v0.4.3, so trailing newlines that had been present in v0.4.1 were removed, leading to a slight accuracy shift in MMLU tasks. After patching, accuracy matched across both versions.
Further Investigation:
Few-shot settings showed matching results after reordering and newline patches were applied, ensuring consistency between v0.4.1 and v0.4.3 in few-shot evaluations. When multiplying the accuracies by the number of samples in the arc_challenge_FR test split (1172 samples), the higher accuracy resulted in 472 correct answers, while the lower accuracy yielded 471 correct answers. This marginal difference (0.085%) may be explained by numerical differences between log-probabilities during batching and ultimately the discontinuous nature of the Accuracy metric.
Conclusion:
The differences in batching and Jinja2 templating in v0.4.3 explain the accuracy variations observed in this experiment. The patched version of v0.4.3 aligns with v0.4.1 when these issues are addressed, suggesting that changes in batching or task description formatting may lead to non-deterministic results across different hardware or LM Eval Harness versions. While the accuracy variations were small (less than 1%), the discrepancies were linked to batching behavior and templating in v0.4.3. Ensuring consistency across these factors is crucial for accurate model evaluations, especially when comparing models across different versions. It should be a development goal for LLM evaluation frameworks that, except for padding, the inputs actually passed to the model under test should not vary depending on batch size, degree of data or model parallelism or other tasks evaluated at the same time.
B.3 Downstream results tables
Tables LABEL:tab:acc_std_21lang_1 and LABEL:tab:acc_std_21lang_2 contain model- and task-wise arithmetic means and population standard deviations across the EU21 languages, and the average of these means. Tables LABEL:tab:acc_std_lang_group_germanic_1, LABEL:tab:acc_std_lang_group_germanic_2, LABEL:tab:acc_std_lang_group_romance_1, LABEL:tab:acc_std_lang_group_romance_2, LABEL:tab:acc_std_lang_group_slavic_1, and LABEL:tab:acc_std_lang_group_slavic_2 contain the same information across the languages of the Germanic, Romance and Slavic language families instead of the EU21 languages. Table 20 lists the differences of the per-task mean accuracies of the high-resource languages minus the corresponding mean accuracies of the medium-resource languages according to table 1.